OpenStudy (anonymous):

1. Tony and Belinda have a combined age of 56. Belinda is 8 more than twice Tony’s age. How old is each?

OpenStudy (mertsj):

t+b=56 b=2t+8

OpenStudy (anonymous):

what do u do next?

OpenStudy (mertsj):

Solve that system

OpenStudy (mertsj):

Use substitution by replacing b in the first equation with 2t+8 from the second equation.

OpenStudy (anonymous):

t+b=56, <equation 1 b-8=2t <equation 2 so now solve - i will solve by substitution t=56-b sub: b-8=2t b-8=2(56-b) b-8=112-2b b+2b=120 3b=120 b=40, so belinda is 40, sub to find tony's age t+b=56 t=56-40 t=16 therefore, tony is 16 and belinda is 40. i think this is the answer

OpenStudy (anonymous):

@dg98 Thank You!!